OpenStudy (anonymous):

If the equation is a parabola state the vertex, focal point (focus), directrix and graph .Remember that the directrix is a line and should be written as x = , or y = some number. If the equation is an ellipse, state the center point (h, k), the lengths of the major and minor axis, the vertices, the focal points (foci), the eccentricity and graph. If the equation is a hyperbola, state the center point (h, k), the vertices, the focal points (foci), the eccentricity, the asymptotes and graph.

OpenStudy (anonymous):

1) – 8y = x2 2) y2 + 6y + 6 – 3x = 0 3) 25x2 + 16y2 = 400 4) 4x2 + 9y2 – 16x + 90y + 97 = 0 5) y2